Question 2 A theater group made appearances in two cities. The hotel charge before tax in the second city was $500 lower than in the first. The tax in the first city was 9% , and the tax in the second city was 3% . The total hotel tax paid for the two cities was $465 . How much was the hotel charge in each city before tax?
Answer by mananth(16946)

A theater group made appearances in two cities. The hotel charge before tax in the second city was $500 lower than in the first. The tax in the first city was 9% , and the tax in the second city was 3% . The total hotel tax paid for the two cities was $465 . How much was the hotel charge in each city before tax?Hotel charge before tax in first city be x ; tax 9%
other city (x-500) tax 3%
Total tax 465
9%x+(x+500)*3% = 465
multiply equation by 100
9x+3(x+500)=46500
12x+1500 =46500
12x = 45000
x= 3750
Hotel charge in City I $3750
subtract 500 to get the other city hotel charge